OpenStudy (anonymous):

write each percent as a decimal and as a ratio 1. 63% 2.8.9% 3.225% 4.19 4/5

OpenStudy (anonymous):

Per cent means per 100. 1. 63 per 100 = 63 / 100 = .63 = 63:100 2. 8.9 per 100 = 8.9 / 100 = .089 = 89:1000 3. 22.5 per 100 = 22.5 / 100 = .225 = 225:1000 = 9:40 4. 19 4/5% = 19.8% = 19.8 per 100 = 19.8/100 = .198 = 198:1000 = 99:500

OpenStudy (anonymous):

Thank you